ABCD is a square. The length of the diagonal is 10√ ̅2. Find the perimeter of the sqare.
Answer by mananth(16946)

ABCD is a square. The length of the diagonal is 10√ ̅2. Find the perimeter of the sqare.

The area of square = 1/2 * (product of diagonals)
Area = 1/2 * ( 10sqrt(2) * 10sqrt(2)}}}
= 1/2 * 200
=100
area = 100
so side = sqrt(100)
= 10
Perimeter = 4*10 = 40